40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3956     Entry time: Fri Nov 19 16:13:09 2010
Author: Jenne 
Type: Update 
Category: SUS 
Subject: ETMU05: magnets glued to optic 

[Jenne, Suresh]

Suresh and I glued the intact-from-the-first-round magnets to ETMU05.  I accidentally got too much glue on one of the dumbbells (the glue was connecting the dumbbell to the gripper - bad news if we let that dry), and while I was cleaning it, the magnet broke off.  So I used one of the ones that Suresh had re-glued last night, and he is putting that one back together after some cleaning. 

To set the fixture, Suresh had the great idea of using small pieces of foil underneath the teflon pads to set the height of the optic in the fixture.  The optic still rests on the teflon pads, but with the foil we have finer control over how the optic sits.  Neat.  Since both ETMs are the same, we shouldn't have to do any more adjustment for the other ETM.

The updated Status Table:

StatusTable.png

ELOG V3.1.3-